OpenStudy (anonymous):

Which angle corresponds to 1/6 of a circle? 1) pie/4 2) pie/6 3) pie/3 4) 180

OpenStudy (freckles):

angle you seek is 1/6 * 360 deg this will give you the angle in degrees 1/6 * 360*pi/180 will give the angle in radians

OpenStudy (freckles):

and i think you mean pi not pie i like to pie but eating pi is impossible because pi is a number :p

OpenStudy (freckles):

i like to eat pie*

OpenStudy (anonymous):

Lmfaooooooo, o the answer would be 180? sorry im helping my daughter, i have no idea what im doing :)

OpenStudy (freckles):

well it wouldn't be 180

OpenStudy (freckles):

\[\frac{1}{6} \cdot 360^o=(\frac{360}{6})^o \text{ is answer \in degrees } \\ \text{ to convert deg to rad} \\ \text{ multiply by } \frac{ \pi}{180^o} radians \]

OpenStudy (freckles):

so that means you need to do \[\frac{\pi}{180^o} \cdot \frac{360}{6}^o radians \]

OpenStudy (freckles):

leave the pi there just worry about computing 360/(6*180)

OpenStudy (freckles):

then bring down the pi

OpenStudy (freckles):

\[\text{ that is } \\ \frac{360^o}{180^o \cdot 6} \pi \text{ radians }\]

OpenStudy (freckles):

the deg units will cancel

OpenStudy (freckles):

360/180=?

OpenStudy (freckles):

here I will try to look even easier to compute \[\text{ that is } \\ \frac{360^o}{180^o \cdot 6} \pi \text{ radians } \\ = \frac{360/180}{6} \pi \text{ radians }\]

OpenStudy (anonymous):

2

OpenStudy (anonymous):

sorry i was finding a calculator

OpenStudy (freckles):

so that means we have \[\text{ that is } \\ \frac{360^o}{180^o \cdot 6} \pi \text{ radians } \\ = \frac{360/180}{6} \pi \text{ radians } \\ =\frac{2 }{6} \pi \text{ radians }\]

OpenStudy (freckles):

now just reduce the fraction 2/6

OpenStudy (anonymous):

3

OpenStudy (freckles):

and after that don't forget to bring down the pi well it would be 1/3 since 6 is on bottom 6/2=3 so 2/6=1/3 so 2pi/6=1pi/3 or just pi/3

OpenStudy (anonymous):

thank you Freckles!!!!

OpenStudy (freckles):

np